A quality control expert at LIFE batteries wants to test their new batteries. The design engineer claims they have a variance of 5929 with a mean life of 941 minutes. If the claim is true, in a sample of 109 batteries, what is the probability that the mean battery life would be greater than 948.8 minutes? Round your answer to four decimal places.
Answer #1

Solution :

Given that,

mean = \mu = 941

variance = \sigma 2 =  5929

standard deviation = \sigma =77

n = 109

\mu\bar x = 941

\sigma\bar x = \sigma / \sqrt n = 77 \sqrt 109 = 7.3751

P (\bar x > 948.8 )

= 1 - P (\bar x <948.8 )

= 1 - P ( \bar x -  \mu\bar x/\sigma\bar x ) < ( 948.8 - 941 / 7.3751)

= 1 - P ( z < 7.8 / 7.3751 )

= 1 - P ( z <1.06 )

Using z table

= 1 - 0.8554

= 0.1446

Probability = 0.1446